# SCott: Accelerating Diffusion Models with Stochastic Consistency Distillation

### v1

@@ Abstract
The iterative sampling procedure employed by diffusion models (DMs) often leads to significant inference latency. To address this, we propose Stochastic Consistency Distillation (SCott) to enable accelerated text-to-image generation, where high-quality generations can be achieved with just $1-2$ sampling steps, and further improvements can be obtained by adding additional steps. In contrast to vanilla consistency distillation (CD) which distills the ordinary differential equation solvers-based sampling process of a pre-trained teacher model into a student, SCott explores the possibility and validates the efficacy of integrating stochastic differential equation (SDE) solvers into CD to fully unleash the potential of the teacher. SCott is augmented with elaborate strategies to control the noise strength and sampling process of the SDE solver. An adversarial loss is further incorporated to strengthen the sample quality with rare sampling steps. Empirically, on the MSCOCO-2017 5K dataset with a Stable Diffusion-V1.5 teacher, SCott achieves an FID (Fréchet Inception Distance) of 22.1, surpassing that (23.4) of the 1-step InstaFlow (Liu et al., 2023) and matching that of 4-step UFOGen (Xu et al., 2023b). Moreover, SCott can yield more diverse samples than other consistency models for high-resolution image generation (Luo et al., 2023a), with up to $16\%$ improvement in a qualified metric. The code and checkpoints are coming soon.
1 Introduction
@@
Diffusion models (DMs) (Ho et al., 2020; Sohl-Dickstein et al., 2015; Song et al., 2020c) have emerged as a pivotal component in the realm of generative modeling, facilitating notable progress in domains including image generation (Ramesh et al., 2022; Rombach et al., 2022), video synthesis (Blattmann et al., 2023; Ho et al., 2022), and beyond. In particular, latent diffusion models (LDMs), such as Stable Diffusion (Rombach et al., 2022), have exhibited exceptional capabilities for high-resolution text-to-image synthesis and are acting as fundamental building components for a wide spectrum of downstream applications (Gal et al., 2022; Chen et al., 2023; Mou et al., 2023).
@@
However, it is widely recognized that DMs’ iterative reverse sampling process leads to slow inference. One remediation is improving the solvers used for discretizing the reverse process (Song et al., 2020a; Lu et al., 2022a, b), but it is still hard for them to generate within a limited number of steps (e.g., $5$) due to the inevitable discretization errors. Another strategy involves distilling an ordinary differential equation (ODE) based generation process of a pre-trained DM into a shorter one (Salimans & Ho, 2022; Meng et al., 2023), but the distillation cost is routinely high, and the sample quality degrades compared with the full-step teacher. Alternatively, consistency distillation (CD) trains consistency models (CMs) to fit the consistency mappings characterized by the diffusion ODE for few-step generation (Song et al., 2023), and latent consistency model (LCM) (Luo et al., 2023a) applies CD to the latent space of a pre-trained autoencoder to enable high-resolution image generation, but its sample quality is poor within $1-4$ sampling steps. Recently, InstaFlow (Liu et al., 2023), UFOGen (Xu et al., 2023b), and ADD (Sauer et al., 2023b) have succeeded in faithfully generating high-resolution images in just $1-2$ steps, but they share the limitation of failing to trade additional sampling steps for improved outcomes.
@@
The quest for a large-scale text-to-image DM capable of generating high-quality outputs with minimal sampling steps (e.g., $2$) while also benefiting from the advantages of additional steps remains an unresolved and crucial challenge. We motivate our solution with observations that CMs enjoy the cost-quality trade-off by alternating denoising and noise injection at inference time, but current CD approaches have not fully unleashed the potential of the teacher, considering that for a well-trained DM, ODE-based solvers usually underperform stochastic differential equation (SDE) ones with adequate sampling steps  (Xu et al., 2023a; Karras et al., 2022; Gonzalez et al., 2023). This is verified by results in Table 1 as well. Empowered by these, we aim to develop a new approach, named Stochastic Consistency Distillation (SCott), to combine CD with SDE solvers to accelerate the sampling of high-resolution images from LDMs.
@@
Figure 1: $512\times 512$ resolution images generated by SCott using 2 sampling steps. SCott is trained from Realistic-Vision-v51.
@@
By (Song et al., 2023), CMs are originally defined and learned based on ODE solvers. Naturally, a straightforward adoption of regular SDE solvers suffers from low training stability and poor convergence. To address this, we empirically identify several critical factors that render SCott workable. On one hand, we find it necessary to keep the injected noise in SDE solvers at a moderate intensity to stabilize training while enjoying benefits from the stochasticity inherent in SDE. On the other hand, we find it vital to extend the one-step sampling strategy employed in vanilla CD to a multi-step one to further diminish the discretization errors. The multi-step SDE solver also aids in correcting the accumulated errors in the sampling path, thanks to the injection of random noise (Xu et al., 2023a). With these, we obtain a stronger and more versatile teacher for CD.
We introduce a simple Gaussian formula to the student model and the loss function of CD to accommodate the uncertainty in SDE. Surprisingly, such uncertainty can lead to diverse generations for the distilled student model (see Section D.4). From another perspective, the SDE solver implicitly makes stochastic data augmentation for CD, which aids in mode coverage. To encourage the student to yield a good enough approximate sample of real data, an adversarial learning loss is incorporated to correct student output, further boosting the sample quality at $1-4$ sampling steps.
@@
Extensive experiments validate the efficacy of SCott in generating high-quality images with conspicuous details. On MSCOCO-2017 5K validation dataset with a Stable Diffusion-V1.5 (SD1.5) (Rombach et al., 2022) teacher, our 2-step method achieves an FID (Heusel et al., 2017) of 22.1, surpassing the previous state-of-the-art, e.g., 2-step LCM (Luo et al., 2023a) (30.4) and 1-step InstaFlow (Liu et al., 2023) (23.4) and UFOGen (Xu et al., 2023b) (22.5). Besides, SCott can smoothly improve the sample quality with increasing sampling steps. At a 4-step inference, SCott consistently outperforms LCM, InstaFlow, and UFOGen with $1-4$ steps in terms of both FID and CLIP Score (Hessel et al., 2021). On MJHQ-5K validation dataset with Realistic-Vision-v51 (RV5.1) [^1] https://huggingface.co/stablediffusionapi/realistic-vision-v51. teacher, 2-step SCott surpasses LCM by a remarkable margin (24.9 v.s. 37.2 in FID, 0.301 v.s. 0.296 in CLIP Score). For the Coverage metric measuring sample diversity (Naeem et al., 2020), 2-step SCott obtains 0.1287 and 0.1778 gains over 2-step LCM on MSCOCO-2017 5K and MJHQ-5K, respectively.
@@
  • We propose SCott, which accelerates diffusion models to generate high-quality outputs within $1-4$ steps while maintaining the capability for further improvement with additional steps.
@@
  • We explore crucial factors that render SCott workable and integrate adversarial learning objectives into SCott to further improve the few-step sample quality.
@@
  • SCott achieves 1) a state-of-the-art FID score of 22.1 in 2 steps, surpassing competing baselines such as 1-step InstaFlow (23.4) and 2-step LCM (30.4), and matching 4-step UFOGen (22.1), and 2) much higher sample diversity, reflected by the Coverage metric, than LCM (0.9169 v.s. 0.7882).
2 Related Works
2.1 Diffusion Models
@@
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020c, 2021; Song & Ermon, 2020; Karras et al., 2022; Dhariwal & Nichol, 2021) progressively perturb data to Gaussian noise and are trained to denoise the noise-corrupted data. During inference, diffusion models create samples from Gaussian distribution by reversing the noising process. They have achieved unprecedented success in text-to-image generation (Saharia et al., 2022; Ramesh et al., 2022), image inpainting (Lugmayr et al., 2022), and image editing (Meng et al., 2021). To effectively improve the sample quality of conditioned diffusion models, classifier-free guidance (CFG) (Ho & Salimans, 2022) technique is proposed without extra network training.
2.2 Diffusion Acceleration
@@
One of the primary challenges that hinder the practical adoption of diffusion models is the issue of sampling speed due to multiple iterations. Several approaches have been proposed to enhance the sampling efficiency of diffusion models. One type of methods concentrate on training-free numerical solvers (Song et al., 2020b; Lu et al., 2022a, b), such as Denoising Diffusion Implicit Model (DDIM) (Song et al., 2020b) and DPM++ (Lu et al., 2022b). Some researchers explore the approaches of knowledge distillation to compress sampling steps. Progressive Distillation (PD) (Salimans & Ho, 2022) and Classifire-aware Distillation (CAD) (Meng et al., 2023) are designed to reduce sampling steps to below 10 steps via multi-stage distillation. LCM (Luo et al., 2023a) extends CM (Song et al., 2023) to text-to-image generation with few-step inference. However, these methods synthesize blurry samples in four steps.
InstaFlow (Liu et al., 2023) proposes one-step sampling model for text-to-image generation by combining DMs and Rectified Flow (Liu et al., 2022). Nevertheless, InstaFlow is unable to extend its sampler to multiple steps, and the synthesized images are not satisfactory enough with a single step only. Consistency trajectory models (CTM) (Kim et al., 2023) and Diff-Instruct (Luo et al., 2023b) distill a pre-trained DM into a single-step generator, but their performance on large-scale text-to-image generation is unclear. Motivated by CMs and knowledge distillation, we propose Stochastic Consistency Distillation to generate high-quality images in few-step inference. Our method is able to not only produce high-quality samples with a 2-step sampler, but also improve model performance with increasing steps. Furthermore, due to the introduced stochasticity, our method exhibits better sample diversity.
2.3 Diffusion GANs
@@
Denoising Diffusion GAN (DDGAN) (Xiao et al., 2021) first introduces an adversarial objective to model the non-Gaussian distribution of the denoising process with large step sizes. With GANs as a core technique, UFOGen (Xu et al., 2023b) proposes a one-step diffusion GAN for text-to-image generation. However, the image quality can not be significantly improved or even gets worse when increasing sampling steps for UFOGen. Adversarial Diffusion Distillation (Sauer et al., 2023b) distills pre-trained SD models by GANs and score distillation (Poole et al., 2022), achieving one-step generation. However, the training source is not described in the paper, e.g., training time and training data, and hence we can not make a fair comparison with it. Different from UFOGen and ADD which adopt adversarial learning as their core component, we propose to leverage adversarial loss to assist SCott in producing high-quality images at few-step sampling. Consequently, our method inherits the property of CM which heightens text-to-image alignment and image sharpness with increasing sampling steps, and the image quality is further enhanced by GAN at few-step sampling.
3 Preliminary
@@
Let ${\bm{x}}\in\mathbb{R}^{k}$ denotes a sample from the data distribution $p_{\rm{data}}({\bm{x}})$ and $p({\bm{z}}_{t})=\int p_{\rm{data}}({\bm{x}})\mathcal{N}({\bm{z}}_{t};\alpha_{t}{\bm{x}},\sigma_{t}^{2}\mathbf{I})\mathrm{d}{\bm{x}},\forall t\in[0,T]$ the marginal distribution specified by the forward diffusion process. $\alpha_{t}$ and $\sigma_{t}$ are positive real-valued functions defining the diffusion schedule so that $p({\bm{z}}_{0})=p_{\rm{data}}({\bm{x}})$ and $p({\bm{z}}_{T})\approx\mathcal{N}({\bm{z}}_{T};\mathbf{0},\tilde{\sigma}^{2}\mathbf{I})$ for some $\tilde{\sigma}$. A DM ${\bm{\epsilon}}_{\theta}(\cdot,t):\mathbb{R}^{k}\to\mathbb{R}^{k}$ is trained under score matching principles (Vincent, 2011; Song & Ermon, 2019; Ho et al., 2020) for reversing the diffusion process. [^2] The paper focuses on the $\epsilon$-prediction type of DMs. Other parameterizations are equivalent in theory.
@@
According to the SDE/ODE explanation of the reverse process of DMs (Song et al., 2020c), we can obtain an approximate sample of $p_{\rm{data}}({\bm{x}})$ by drawing a Gaussian noise ${\bm{z}}_{T}\sim p({\bm{z}}_{T})$ and then invoking a numerical SDE/ODE solver to discretize the reverse process. Let $\hat{{\bm{z}}}_{t}$ denotes the solution at timestep $t$, originating from ${\bm{z}}_{T}$ based on a solver and model ${\bm{\epsilon}}_{\theta}(\cdot,t)$, and then $\hat{{\bm{z}}}_{0}$ represents the sampled data.
3.1 Consistency Distillation
@@
The mentioned solving process usually hinges on tens or hundreds of steps, causing significant inference overheads for practical application. A promising solution is to perform consistency distillation (CD) (Song et al., 2023; Song & Dhariwal, 2023) of DMs, yielding a student model ${\bm{f}}_{\bm{\theta}}(\cdot,t):\mathbb{R}^{k}\to\mathbb{R}^{k}$ which enjoys a shortened sampling procedure.
Concretely, CD defines a novel, continuous parameterization of ${\bm{f}}_{\bm{\theta}}({\bm{z}},t)$ so that ${\bm{f}}_{\bm{\theta}}({\bm{z}},t)={\bm{z}}$ when $t$ is tiny, and minimizes the following loss for training
$\small\min_{{\bm{\theta}}}\mathcal{L}_{CD}({\bm{\theta}})=\mathbb{E}_{n,{\bm{z}}_{t_{n}}}\Big{[}\lambda(t_{n})\big{\|}{\bm{f}}_{\bm{\theta}}({\bm{z}}_{t_{n}},{t_{n}})-{\bm{f}}_{{{\bm{\theta}}}^{-}}(\hat{{\bm{z}}}_{t_{m}},t_{m})\big{\|}_{2}^{2}\Big{]}$ (1)
@@
where $m\in\{1,\dots,n-1\}$ is a hyper-parameter (Song et al., 2023; Luo et al., 2023a), $\lambda(\cdot)$ refers to another positive weighting function, ${{\bm{\theta}}}^{-}$ denotes the exponential moving average (EMA) of ${{\bm{\theta}}}$, and
@@
$\begin{split}\small\hat{{\bm{z}}}_{t_{m}}=&\frac{\sqrt{\alpha_{t_{m}}}}{\sqrt{\alpha_{t_{n}}}}({\bm{z}}_{t_{n}}-\sqrt{1-\alpha_{t_{n}}-\sigma_{t_{n}}}\cdot{\bm{\epsilon}}_{\theta}({\bm{z}}_{t_{n}},t_{n}))\\ &\quad\quad\quad\quad+\sqrt{1-\alpha_{t_{m}}}\cdot{\bm{\epsilon}}_{\theta}({\bm{z}}_{t_{n}},t_{n})+\sigma_{t_{n}}\epsilon\end{split}$ (2)
I.e., $\hat{{\bm{z}}}_{t_{m}}$ represents an intermediate state, starting from ${\bm{z}}_{t_{n}}$ and obtained with the teacher model ${\bm{\epsilon}}_{\theta}(\cdot,t)$ and a one-step DDIM solver (Song et al., 2020b). CD typically sets $\sigma_{t_{n}}=0$ to keep $\hat{{\bm{z}}}_{t_{m}}$ and ${\bm{z}}_{t_{n}}$ on the same ODE trajectory.
@@
Although we confine the distance measure in Equation 1 to the squared $\ell_{2}$ distance, the $\ell_{1}$ distance also applies here. We do not consider the Learned Perceptual Image Patch Similarity (LPIPS, Zhang et al., 2018) because it can lead to inflated FID scores (Song & Dhariwal, 2023). The trained ${\bm{f}}_{\bm{\theta}}$ allows for one-step or multi-step sampling for generating new data (Song et al., 2023).
@@
Limitation. A primary limitation of CD from the viewpoint of distillation is that it has not fully unleashed the potential of the model ${\bm{\epsilon}}_{\theta}(\cdot,t)$: it utilizes one-step DDIM to sample a preceding state of the current one to serve as the teacher for distillation. Yet, for a well-trained DM, ODE-based solvers unusually underperform SDE ones with adequate sampling steps (Gonzalez et al., 2023). Naturally, we ask if we can build an SDE-based teacher for improved CD.
3.2 SDE Solvers for Diffusion Models
@@
The SDE formulation of the reverse-time diffusion process (Song et al., 2020c) takes the form of
@@
$\scriptsize\begin{split}\mathrm{d}{{\bm{x}}_{t}}=\underbrace{[f_{t}{\bm{x}}_{t}+\frac{g^{2}_{t}}{2\sigma_{t}}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)]\mathrm{d}t}_{\text{Probabilistic ODE}}+\underbrace{\frac{g^{2}_{t}}{2}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)\mathrm{d}t+g_{t}\mathrm{d}\bar{{\bm{w}}}_{t}}_{\text{Langevin process}}\end{split}$ (3)

### v4

@@ Abstract
The iterative sampling procedure employed by diffusion models (DMs) often leads to significant inference latency. To address this, we propose Stochastic Consistency Distillation (SCott) to enable accelerated text-to-image generation, where high-quality and diverse generations can be achieved within just 2-4 sampling steps. In contrast to vanilla consistency distillation (CD) which distills the ordinary differential equation solvers-based sampling process of a pre-trained teacher model into a student, SCott explores the possibility and validates the efficacy of integrating stochastic differential equation (SDE) solvers into CD to fully unleash the potential of the teacher. SCott is augmented with elaborate strategies to control the noise strength and sampling process of the SDE solver. An adversarial loss is further incorporated to strengthen the consistency constraints in rare sampling steps. Empirically, on the MSCOCO-2017 5K dataset with a Stable Diffusion-V1.5 teacher, SCott achieves an FID of 21.9 with 2 sampling steps, surpassing that of the 1-step InstaFlow (23.4) and the 4-step UFOGen (22.1). Moreover, SCott can yield more diverse samples than other consistency models for high-resolution image generation, with up to $16\%$ improvement in a qualified metric.
Introduction
@@
Diffusion models (DMs) (Ho, Jain, and Abbeel 2020; Sohl-Dickstein et al. 2015; Song et al. 2020) have emerged as a pivotal component in the realm of generative modeling, facilitating notable progress in domains including image generation (Ramesh et al. 2022; Rombach et al. 2022), video synthesis (Blattmann et al. 2023; Ho et al. 2022), and beyond. In particular, latent diffusion models (LDMs), such as Stable Diffusion (Rombach et al. 2022), have exhibited exceptional capabilities for high-resolution text-to-image synthesis and are acting as fundamental building components for a wide spectrum of downstream applications (Gal et al. 2022; Chen et al. 2023b; Mou et al. 2023).
@@
However, it is widely recognized that DMs’ iterative reverse sampling process leads to slow inference. One remediation is improving the solvers used for discretizing the reverse process (Song, Meng, and Ermon 2020; Lu et al. 2022a, b), but it is still hard for them to generate within a limited number of steps (e.g., $5$) due to the inevitable discretization errors. Another strategy involves distilling an ordinary differential equation (ODE) based generation process of a pre-trained DM into a shorter one (Salimans and Ho 2022; Meng et al. 2023), but the cost of such progressive distillation is routinely high. Alternatively, consistency distillation (CD) trains unified consistency models (CMs) to fit the consistency mappings characterized by the diffusion ODE for few-step generation (Song et al. 2023). Latent consistency model (LCM) (Luo et al. 2023a) further applies CD to the latent space of a pre-trained autoencoder to enable high-resolution image generation. However, its sample quality is poor within 2 sampling steps. Recently, InstaFlow (Liu et al. 2023), UFOGen (Xu et al. 2024), and ADD (Sauer et al. 2023b) have succeeded in faithfully generating high-resolution images in just 1-2 steps, but they share the limitation of failing to trade additional sampling steps for improved outcomes.
@@
One-step samplers are hard to generate satisfactory outputs, and numerous works (Luo et al. 2023a; Salimans and Ho 2022; Meng et al. 2023) improve their one-step performance via additional steps with acceptable cost (e.g., 4 steps). Our objective is to strike a unified model capable of generating high-quality outputs with 2-4 steps. This is based on the fact that over-emphasizing the one-step generation capacity would unavoidably bias the DM, hence weakening the multi-step generation capacity. Sampling with 2-4 steps does not substantially increase the practical cost compared to 1 step but is likely to improve the upper bound of the sampling quality significantly. We base our solution on CMs because they enjoy the cost-quality trade-off by alternating denoising and noise injection at inference time. Yet, current CD approaches have not fully unleashed the potential of the teacher, considering that for a well-trained DM, ODE-based solvers usually underperform stochastic differential equation (SDE) ones with adequate sampling steps  (Xu et al. 2023; Karras et al. 2022; Gonzalez et al. 2023). This is verified by results in Table 1. Empowered by these, we aim to develop Stochastic Consistency Distillation (SCott), to combine CD with SDE solvers to accelerate the sampling of high-resolution images.
@@
Figure 1: $512\times 512$ resolution images generated by SCott using 2 sampling steps. SCott is trained based on Realistic-Vision-v51.
@@
By (Song et al. 2023), CMs are originally defined and learned based on ODE solvers. Naturally, a straightforward adoption of regular SDE solvers suffers from low training stability and poor convergence. To address this, we provide a theoretical justification for using SDE solvers for CD and empirically identify several critical factors that render SCott workable. On one hand, we find it necessary to keep the injected noise in SDE solvers at a moderate intensity to stabilize training while enjoying benefits from the stochasticity inherent in SDE. On the other hand, we find it vital to extend the one-step sampling strategy employed in vanilla CD to a multi-step one to further diminish the discretization errors. The multi-step SDE solver also aids in correcting the accumulated errors in the sampling path, thanks to the injection of random noise (Xu et al. 2023). With these, we obtain a stronger and more versatile teacher for CD.
@@
Surprisingly, we find the uncertainty in SDE can lead to diverse generations for the distilled student model (see Appendix 5). From another perspective, the SDE solver implicitly makes stochastic data augmentation for CD, which aids in mode coverage. To mitigate the imperfect consistency constraints provided by L2 loss in CD, an adversarial learning loss is incorporated to correct student output, further boosting the sample quality at 1-4 sampling steps. Extensive experiments validate the efficacy of SCott in generating high-quality images with conspicuous details. On MSCOCO-2017 5K validation dataset with a Stable Diffusion-V1.5 (SD1.5) (Rombach et al. 2022) teacher, our 2-step method achieves an FID (Heusel et al. 2017) of 21.9, surpassing the previous state-of-the-art, e.g., 2-step LCM (Luo et al. 2023a) (30.4) and 1-step InstaFlow (Liu et al. 2023) (23.4) and UFOGen (Xu et al. 2024) (22.5). Besides, SCott can smoothly improve the sample quality with increasing sampling steps. At a 4-step inference, SCott consistently outperforms LCM, InstaFlow, and UFOGen with $2-4$ steps in terms of both FID and CLIP Score (Hessel et al. 2021). On MJHQ-5K validation dataset with Realistic-Vision-v51 (RV5.1) [^1] https://huggingface.co/stablediffusionapi/realistic-vision-v51. teacher, 2-step SCott surpasses LCM by a remarkable margin (24.9 v.s. 37.2 in FID, 0.301 v.s. 0.296 in CLIP Score). For the Coverage metric measuring sample diversity (Naeem et al. 2020), 2-step SCott obtains 0.1232 and 0.1778 gains over 2-step LCM on MSCOCO-2017 5K and MJHQ-5K, respectively.
@@
  • We propose SCott, which accelerates diffusion models to generate high-quality outputs with $1-2$ steps while maintaining the capability for further improvement via more inference time within $4$ steps.
@@
  • We provide theoretical convergence analysis for SCott and explore crucial factors that render SCott workable. Furthermore, we integrate adversarial learning objectives into SCott to improve the few-step sample quality.
@@
  • SCott achieves 1) a state-of-the-art FID of 21.9 in 2 steps, surpassing competing baselines such as 1-step InstaFlow (23.4), 2-step LCM (30.4), and 4-step UFOGen (22.1), and 2) much higher sample diversity, reflected by the Coverage metric, than LCM (0.9114 v.s. 0.7882).
Related Works
Diffusion Models
@@
Diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020; Song et al. 2020, 2021; Song and Ermon 2020; Karras et al. 2022; Dhariwal and Nichol 2021) progressively perturb data to Gaussian noise and are trained to denoise the noise-corrupted data. During inference, diffusion models create samples from Gaussian distribution by reversing the noising process. They have achieved unprecedented success in text-to-image generation (Saharia et al. 2022; Ramesh et al. 2022), image inpainting (Lugmayr et al. 2022), and image editing (Meng et al. 2021; Chen et al. 2023a). To effectively improve the sample quality of conditioned diffusion models, classifier-free guidance (CFG) (Ho and Salimans 2022) technique is proposed without extra network training.
Diffusion Acceleration
@@
One of the primary challenges that hinder the practical adoption of diffusion models is the issue of sampling speed due to multiple iterations. Several approaches have been proposed to enhance the sampling efficiency of diffusion models. One type of methods concentrate on training-free numerical solvers (Song, Meng, and Ermon 2020; Lu et al. 2022a, b), such as Denoising Diffusion Implicit Model (DDIM) (Song, Meng, and Ermon 2020) and DPM++ (Lu et al. 2022b). Some researchers explore the approaches of knowledge distillation to compress sampling steps. Progressive Distillation (PD) (Salimans and Ho 2022) and Classifire-aware Distillation (CAD) (Meng et al. 2023) are designed to reduce sampling steps to below 10 steps via multi-stage step distillation. LCM (Luo et al. 2023a) extends CM (Song et al. 2023) to text-to-image generation. However, these methods synthesize blurry samples below four steps. Recently, InstaFlow (Liu et al. 2023), SwiftBrush (Nguyen and Tran 2024), and DMD (Yin et al. 2024) achieve one-step generation in high-resolution text-to-image generation faithfully. InstaFlow proposes a one-step sampling model for text-to-image generation by combining DMs and Rectified Flow (Liu, Gong, and Liu 2022). SwiftBrush adopts variational score distillation (VSD) (Wang et al. 2024) to distill a one-step student. DMD employs distribution matching distillation to enhance the realism of the one-step generator. Nevertheless, they are unable to extend their sampler to multiple steps, and the synthesized images are not satisfactory enough with a single step only. Consistency trajectory models (CTM) (Kim et al. 2023) and Diff-Instruct (Luo et al. 2023b) distill a pre-trained DM into a single-step generator, but their performance on large-scale text-to-image generation is unclear. Motivated by CMs and knowledge distillation, we propose Stochastic Consistency Distillation to generate high-quality images within few steps. Our method is able to not only produce high-quality samples with a 2-step sampler but also improve model performance with increasing steps. Furthermore, due to the introduced stochasticity in SDE solvers, our method exhibits better sample diversity.
Diffusion GANs
@@
With GANs as a core technique, UFOGen (Xu et al. 2024) proposes a one-step diffusion GAN for text-to-image generation. However, the image quality can not be significantly improved or even gets worse when increasing sampling steps for UFOGen. Adversarial Diffusion Distillation (Sauer et al. 2023b) distills pre-trained SD models by GANs and score distillation (Poole et al. 2022), achieving one-step generation. However, the training source is not described in the paper, e.g., training time and training data, and hence we can not make a fair comparison with it. Different from UFOGen and ADD which adopt adversarial learning as their core component, we propose to leverage adversarial loss to strengthen consistency constraints in SCott, producing high-quality images at few-step sampling. Consequently, our method inherits the property of CM which heightens text-to-image alignment and image sharpness with increasing sampling steps, and the image quality is further enhanced by GAN at few-step inference.
Preliminary
@@
Let ${\bm{x}}\in\mathbb{R}^{k}$ denotes a sample from the data distribution $p_{\rm{data}}({\bm{x}})$ and $p({\bm{z}}_{t})=\int p_{\rm{data}}({\bm{x}})\mathcal{N}({\bm{z}}_{t};\alpha_{t}{\bm{x}},\sigma_{t}^{2}\mathbf{I})\mathrm{d}{\bm{x}},\forall t\in[0,T]$ the marginal distribution specified by the forward diffusion process. $\alpha_{t}$ and $\sigma_{t}$ are positive real-valued functions defining the diffusion schedule so that $p({\bm{z}}_{0})=p_{\rm{data}}({\bm{x}})$ and $p({\bm{z}}_{T})\approx\mathcal{N}({\bm{z}}_{T};\mathbf{0},\tilde{\sigma}^{2}\mathbf{I})$ for some $\tilde{\sigma}$. A DM ${\bm{\epsilon}}_{\theta}(\cdot,t):\mathbb{R}^{k}\to\mathbb{R}^{k}$ is trained under score matching principles (Vincent 2011; Song and Ermon 2019; Ho, Jain, and Abbeel 2020) for reversing the diffusion process. [^2] The paper focuses on the $\epsilon$-prediction type of DMs. Other parameterizations are equivalent in theory (Salimans and Ho 2022).
@@
According to the SDE/ODE explanation of the reverse process of DMs (Song et al. 2020), we can obtain an approximate sample of $p_{\rm{data}}({\bm{x}})$ by drawing a Gaussian noise ${\bm{z}}_{T}\sim p({\bm{z}}_{T})$ and then invoking a numerical SDE/ODE solver to discretize the reverse process. Let $\hat{{\bm{z}}}_{t}$ denotes the solution at timestep $t$, originating from ${\bm{z}}_{T}$ based on a solver and model ${\bm{\epsilon}}_{\theta}(\cdot,t)$, and then $\hat{{\bm{z}}}_{0}$ represents the sampled data.
Consistency Distillation
@@
The mentioned solving process usually hinges on tens or hundreds of steps, causing significant inference overheads for practical application. A promising solution is to perform consistency distillation (CD) (Song et al. 2023; Song and Dhariwal 2023) of DMs, yielding a student model ${\bm{f}}_{\bm{\theta}}(\cdot,t):\mathbb{R}^{k}\to\mathbb{R}^{k}$ which enjoys a shortened sampling procedure.
Concretely, CD defines a novel consistency function $\mathbf{f}:(\mathbf{z}_{t},t)\mapsto\mathbf{z}_{\tau}$ for all $t\in[\tau,T]$, where $\tau$ is the boundary time near 0. CD parameterize consistency function with networks ${\bm{f}}_{\bm{\theta}}({\bm{z}}_{t},t)$, and minimizes the following loss for training
$\min_{{\bm{\theta}}}\mathcal{L}_{CD}({\bm{\theta}})=\mathbb{E}_{n,{\bm{z}}_{t_{n}}}\Big{[}\lambda(t_{n})\big{\|}{\bm{f}}_{\bm{\theta}}({\bm{z}}_{t_{n}},{t_{n}})-{\bm{f}}_{{{\bm{\theta}}}^{-}}(\hat{{\bm{z}}}_{t_{m}},t_{m})\big{\|}_{2}^{2}\Big{]},$ (1)
@@
where $t$ in $[\tau,T]$ are uniformly discretized into $N$ time points with $t_{1}=\tau<t_{2}<\cdots<t_{N}=T$ and $m\in\{1,\dots,n-1\}$ is a hyper-parameter  (Song et al. 2023; Luo et al. 2023a), $\lambda(\cdot)$ refers to another positive weighting function, ${{\bm{\theta}}}^{-}$ denotes the exponential moving average (EMA) of ${{\bm{\theta}}}$. The state $\hat{{\bm{z}}}_{t_{m}}$ represents an intermediate state, starting from ${\bm{z}}_{t_{n}}$ and obtained with the teacher model ${\bm{\epsilon}}_{\theta}(\cdot,t)$ and a one-step DDIM solver (Song, Meng, and Ermon 2020), expressed as
@@
$\begin{split}\hat{{\bm{z}}}_{t_{m}}=&\frac{\sqrt{\alpha_{t_{m}}}}{\sqrt{\alpha_{t_{n}}}}({\bm{z}}_{t_{n}}-\sqrt{1-\alpha_{t_{n}}}\cdot{\bm{\epsilon}}_{\theta}({\bm{z}}_{t_{n}},t_{n}))\\ &\quad\quad\quad\quad+\sqrt{1-\alpha_{t_{m}}-\sigma_{t_{n}}^{2}}\cdot{\bm{\epsilon}}_{\theta}({\bm{z}}_{t_{n}},t_{n})+\sigma_{t_{n}}{\bm{\epsilon}}.\end{split}$ (2)
CD typically sets $\sigma_{t_{n}}=0$ to keep $\hat{{\bm{z}}}_{t_{m}}$ and ${\bm{z}}_{t_{n}}$ on the same ODE trajectory.
@@
Although we confine the distance measure in Equation 1 to the squared $\ell_{2}$ distance, the $\ell_{1}$ distance also applies here. We do not consider the Learned Perceptual Image Patch Similarity (LPIPS, Zhang et al. 2018) because it can lead to inflated FID scores (Song and Dhariwal 2023). The trained ${\bm{f}}_{\bm{\theta}}$ allows for one-step or multi-step sampling for generating new data (Song et al. 2023).
@@
Limitation. A primary limitation of CD from the viewpoint of distillation is that it has not fully unleashed the potential of the model ${\bm{\epsilon}}_{\theta}(\cdot,t)$: it utilizes one-step DDIM to sample a preceding state of the current one to serve as the teacher for distillation. Yet, for a well-trained DM, ODE-based solvers unusually underperform SDE ones with adequate sampling steps (Gonzalez et al. 2023). Naturally, we ask if we can build an SDE-based teacher for improved CD.
SDE Solvers for Diffusion Models
@@
The SDE formulation of the reverse-time diffusion process (Song et al. 2020) takes the form of
@@
$\begin{split}\mathrm{d}{{\bm{x}}_{t}}=\underbrace{[f_{t}{\bm{x}}_{t}+\frac{g^{2}_{t}}{2\sigma_{t}}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)]\mathrm{d}t}_{\text{Probabilistic ODE}}+\underbrace{\frac{g^{2}_{t}}{2\sigma_{t}}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)\mathrm{d}t+g_{t}\mathrm{d}\bar{{\bm{w}}}_{t},}_{\text{Langevin process}}\end{split}$ (3)